# Fine-Tuning a Simulation-Driven Estimator

### v2

@@ IV-A Van der Pol Oscillator
$\dot{x}_{1}=x_{2},\qquad\dot{x}_{2}=\mu(1-x_{1}^{2})x_{2}-x_{1}+w_{k},\qquad y=x_{1},$ (20)
with unknown $\mu>0$ and small process noise $w_{k}\sim\mathcal{N}(0,0.02)$. We collect $N=300$ length observations for each $\mu_{i}\sim\mathrm{Unif}[0,2.5]$ to form the offline training set $\mathcal{D}_{\mathrm{tr}}$.
OOD tests. We evaluate at $\mu=3.0$ and $\mu=5.0$, which are outside the pretraining range. Over $n_{\mathrm{mc}}=100$ Monte-Carlo runs, we report mean squared errors (MSEs) and average trajectories of the true system (20) and of each method (Figs. 22). TS-fine has a substantially smaller MSE than TS-pre and outperforms dual-EKF/PEM under poor initializations; PEM with a good initialization can perform well, but is initialization-sensitive, unlike TS-fine.
Figure 1: VDP (OOD $\mu=3.0$): ground-truth vs. mean predictions over $100$ MC runs. The trajectory plots are measurements without noise.
Figure 2: VDP (OOD $\mu=5.0$): ground-truth vs. mean predictions over $100$ MC runs. The trajectory plots are measurements without noise.
IV-A1 Remark on choosing $\alpha$
We sweep $\alpha\in\{0.50,0.30,0.20,0.10,0.05,0.01\}$ on two scenarios ($N=300$, $n_{\mathrm{mc}}=100$ trials): (i) in-distribution (ID, $\mu=1.5$) and (ii) out-of-distribution (OOD, $\mu=4.5$). Empirically, for ID the OOD detection rate drops to 0 by $\alpha=0.05$ and TS-fine offers no benefit in terms of MSE, whereas for OOD the statistical test is positive essentially always and fine-tuning reduces the MSE by orders of magnitude (e.g., $4.70\rightarrow 5\times 10^{-3}$). Thus choosing $\alpha\in[0.05,\,0.10]$ avoids unnecessary fine-tuning on ID while reliably detecting OOD cases.
IV-B Cascaded Water Tanks
@@
OOD tests. We consider $\boldsymbol{\theta}=(1.2,1.2,0.9,1.0)^{\top}$ (OOD $1$) and $\boldsymbol{\theta}=(1.3,1.3,0.6,0.7)^{\top}$ (OOD $2$), both outside the pretraining set. For evaluation we fix the same block-uniform input across runs, simulate with the noise model above, and average over $n_{\mathrm{mc}}=100$ Monte-Carlo (MC) runs. From Fig. 4 and 4, we observe that TS-fine improves upon TS-pre and is competitive with (or better than) dual-EKF/PEM without relying on any initialization. PEM can excel with a good initialization; TS-fine is initialization-free.
Remarks
(i) The accuracy of TS-pre and TS-fine hinge on the specific first and second-stage choices. In particular, the first-stage compressor should be robust to measurement noise, and our current second stage—an MLP (with LayerNorm for tanks)—can be brittle in low SNR settings. Developing more noise-robust alternatives for both stages is left to future work. (ii) Our aim here is to fine-tune TS, not to optimize the underlying first/second-stage architectures for noise robustness. Consequently, in low SNR regimes, the standard baselines may outperform our current setup.
Figure 3: Tanks (OOD 1): average trajectory (ground truth vs. mean predictions over $100$ MC runs. The trajectory plots are measurements without noise.
@@
V Conclusion
In this paper, we have proposed a transfer-learning approach to fine-tune a pretrained, simulation-driven (TS) estimator for out-of-distribution (OOD) scenarios. We first apply a statistical detection test on the compressed features to flag OOD; if the test detects an OOD, we modify only the second stage (a neural network) by freezing the shared trunk and retraining the last, per-parameter layers using a newly synthesized local dataset. Numerical examples demonstrate that the fine-tuned TS delivers improved performance over the pretrained TS and fares well against initialization-dependent baselines such as dual-EKF and PEM. The main limitation is the computational cost of retraining the heads (and generating the synthetic data); in future work we aim to reduce this overhead.

### v3

@@ 5.1 Van der Pol Oscillator
$\dot{x}_{1}=x_{2},\qquad\dot{x}_{2}=\nu(1-x_{1}^{2})x_{2}-x_{1}+w_{k},\qquad y=x_{1},$ (16)
with unknown $\nu>0$ and small process noise $w_{k}\sim\mathcal{N}(0,0.02)$. We collect $N=300$ length observations for each $\nu_{i}\sim\mathrm{Unif}[0,2.5]$ to form the offline training set $\mathcal{D}_{\mathrm{tr}}$.
OOD tests. We evaluate at $\nu=3.0$ (OOD $1$), a mild OOD and $\nu=5.0$ (OOD $2$), which are outside the pretraining range. Over $n_{\mathrm{mc}}=100$ Monte-Carlo (MC) runs, we report mean squared errors (MSEs) for OOD $1$ and average trajectories of the true system (16) and of each method for OOD $2$ (Figs. 1 and Table 1) [^2] We report OOD 1 quantitatively (Table 1) and OOD 2 qualitatively (Figs. 1, 2), since both OOD settings show the same relative gaps, so we present complementary views.. TS-fine has a substantially smaller MSE than TS-pre and outperforms dual-EKF/PEM under poor initializations; PEM with a good initialization can perform well, but is initialization-sensitive, unlike TS-fine.
Figure 1: VDP (OOD 2): ground-truth vs. mean predictions over $100$ MC runs. The trajectory plots are measurements without noise.
Remark 5.2
We set $\alpha\in[0.05,\,0.10]$.This range was selected from a preliminary sensitivity analysis to balance false positives on in-distribution data with reliable OOD detection. Unless otherwise noted, we use $\alpha=0.10$ in all experiments.
5.2 Cascaded Water Tanks
@@
OOD tests. We consider $\boldsymbol{\theta}=(1.2,1.2,0.9,1.0)^{\top}$ (OOD $1$) and $\boldsymbol{\theta}=(1.3,1.3,0.6,0.7)^{\top}$ (OOD $2$), both outside the pretraining set. For evaluation we fix the same block-uniform input across runs, simulate with the noise model above, and average over $n_{\mathrm{mc}}=100$ MC runs. A comparison of the results in terms of MSE and trajectory tracking is presented in Table 1 for OOD $1$ and Fig. 2. From Fig. 2 and Table 1, we observe that TS-fine improves upon TS-pre and is competitive with (or better than) dual-EKF/PEM without relying on any initialization. PEM can excel with a good initialization; TS-fine is initialization-free.
6 Future work
The following are possible directions for future work:
  1. 1.
    The accuracy of TS-pre and TS-fine depends on both stages. Under high process/measurement noise (low SNR), the Stage 1 compressor should yield statistics from which the parameters can be inferred irrespective of the noise level, while Stage 2 (MLP) can become unreliable during training. Our aim here is to fine-tune TS, not to optimize the underlying first/second-stage architectures for noise robustness. Consequently, in low SNR regimes, the standard baselines may outperform our current setup. This can be addressed by (a) designing a noise-robust Stage 1 compressor and (b) strengthening Stage 2 via adversarial training.
  2. 2.
    Fine-tuning adds a computation overhead from three sources: (i) computing $\boldsymbol{J}$ and solving the $d\times d$ system $\boldsymbol{G}=\boldsymbol{J}^{\top}\boldsymbol{J}+\lambda\boldsymbol{I}_{d}$, which has $\mathcal{O}(d^{3})$ time complexity, (ii) generating a synthetic cloud of $M_{\mathrm{FT}}$ parameter samples and their features, and (iii) short, head-only retraining over this cloud. This computation overhead can be overcome via unsupervised fine-tuning that updates only the final-layer weights by directly shrinking the discrepancy between simulated and observed trajectories/features, thereby avoiding large synthetic clouds.
Table 1: OOD 1 results: MSE and avg. computation time (ms).
VDP OOD 1 Tanks OOD 1
Method MSE ms MSE ms
TS-pre $0.3204$ 1.885 4881.885\,488 $0.1189$ 3.994 9493.994\,949
TS-fine 0.000 550 10.000\,550\,1 13 087.631 39413\,087.631\,394 0.070 510.070\,51 50 854.723 69150\,854.723\,691
Dual-EKF (GI) 0.003 2690.003\,269 8.289 9878.289\,987 $6.112$ 18.850 67118.850\,671
Dual-EKF (TI) 0.006 730.006\,73 8.401 9778.401\,977 $8.345$ 19.106 95719.106\,957
PEM (GI) 0.001 1540.001\,154 138.949 625138.949\,625 0.073 670.073\,67 8564.443 6398564.443\,639
PEM (TI) 0.001 1540.001\,154 115.814 819115.814\,819 $0.186$ 8783.367 6528783.367\,652
Dual-EKF (WI) $15.78$ 8.368 5228.368\,522 $8.978$ 19.033 70219.033\,702
PEM (BI) $14.67$ 83.472 52783.472\,527 $0.3924$ 7086.184 8317086.184\,831
Init: GI = Good, TI = TS-init, WI = Worst-case Init, BI = Bad Init.
@@ 6 Future work
7 Conclusion
In this paper, we have proposed a transfer-learning approach to fine-tune a pretrained, simulation-driven (TS) estimator for OOD scenarios. We first apply a statistical detection test on the compressed features to flag OOD; if the test detects an OOD, we modify only the second stage, a neural network, by freezing the shared trunk and retraining the last, per-parameter layers using a newly synthesized local dataset. Numerical examples demonstrate that the fine-tuned TS delivers improved performance over the pretrained TS and fares well against initialization-dependent baselines such as dual-EKF and PEM.